OpenStudy (anonymous):

What is the solution of the system? Use elimination. 3x – 4y = 9 –3x + 2y = 9

OpenStudy (anonymous):

y=-9/2 => x= -9/3 = -3

OpenStudy (radar):

Are these two equations? 3x-4y=9 -3x + 2y = 9

OpenStudy (anonymous):

A. (3, 9) B. (–27, –9) C. (–3, –6) D. (–9, –9)

OpenStudy (radar):

IF so, I will add them to eliminate the x -2y=18 y=-9 Can you take it from there and solve for x?

OpenStudy (radar):

No dialog..........substitute the -9 for y in the 3x-4y=9 3x+36=9 3x=9-36=-27 x=-9 (-9,-9) Did you follow with understanding?